What is the molar solubility of silver chloride in a 0.05 M KNO3 solution, if the Ksp of silver chloride is 1.8*10^-10?
Answer #1

AgCl ------------> Ag+   + Cl-

1 mole of silver chloride produces one mole of silver cation and 1 mole of chloride anion

Considering molar solubility as x

Then Ksp = X*X = X2

Ksp = 1.8*10-10

Therefore X2 = 1.8*10-10

X = phpyURv5E.png1.8*10-10

molar solubility is = 1.34*10-5 M